OpenStudy (anonymous):

Michiko works in a shop that rents cross-country skis. Two clients come in to rent skis. Michiko knows that a person's height is proportional to his or her correct ski length. She figures out that the first client, who is 70 inches tall, needs 210-centimeter skis. The second client is 63 inches tall. What length of skis does the second client need?

OpenStudy (anonymous):

you can set this up as a proportion.... 70/210 = 63/x (70 in is to 210 cm as 63 in is to x cm) now cross multiply (70)(x) = (210)(63) 70x = 13230 (divide by 70 to get x by itself) 70x/70 = 13230/70 x = 189 check... 70/210 = 63/189 (cross multiply) (70)(189) = (210)(63) 13230 = 13230 (correct) ANSWER : The second client needs a ski length of 189 cm.